# InstructTA: Instruction-Tuned Targeted Attack for Large Vision-Language Models

### v2

@@ 4.1 Evaluation Setup
Datasets. The ImageNet-1K [13] comprises 3 subsets, including training, validation and test sets. We randomly select 1,000 images from the validation set to serve as the benign samples ($\boldsymbol{x}$) for the attack. Correspondingly, we sample 1,000 instruction-answer pairs from LLaVA-Instruct-150K [27] as tested instructions and target texts, respectively, since LLaVA-Instruct-150K is a set of multimodal instruction-following data generated by GPT-4. In addition, we use Stable Diffusion ($h_{\xi}$) to generate target images from the target texts, following [46].
@@
Protocols. Following [46], we adopt CLIP-score ($\uparrow$) to evaluate the targeted attack performance, which calculates the semantic similarity between the generated response of victim LVLM and the target text using CLIP text encoders. Additionally, we record the number of successful attacks by using GPT-4 to determine whether two texts have the same meaning. The prompt in this operation is as follows: Determine whether these two texts describe the same objects or things, you only need to answer yes or no: 1. $\boldsymbol{y}^{\prime}$ 2. $\boldsymbol{y}_{t}$ where $\boldsymbol{y}^{\prime}$ denotes the output response of the adversarial example on victim model, and $\boldsymbol{y}_{t}$ is the target response.
@@
Baselines. This paper provides evaluations on three victim LVLMs, including BLIP-2 [25], InstructBLIP [12] and MiniGPT-4 [48]. In detail, we select flan-t5-xxl [^2] https://huggingface.co/google/flan-t5-xxl, vicuna-7b-v1.1333https://huggingface.co/lmsys/vicuna-7b-v1.1 and vicuna-v0-13b [^4] https://huggingface.co/Vision-CAIR/vicuna as LLMs for BLIP-2, InstructBLIP and MiniGPT-4, respectively. Their visual encoders are all ViT-G/14 [42] borrowed from EVA-CLIP [17]. To evaluate the targeted attack performance, InstructTA is compared with existing baselines, i.e., MF-it and MF-ii. They can choose CLIP (ViT-B/32) [34] or the EVA-CLIP (ViT-G/14) as surrogate models for targeted attacks.
@@
Implementation Details. All images are resized to $224\times 224$ before feeding in substitute models and LVLMs. For InstructTA, we choose the visual encoder and Q-Former from InstructBLIP as the surrogate model. For PGD, $\eta$ and $T$ are set to 1 and 100, respectively. The maximum perturbation magnitude $\epsilon$ is set to $8$. We implement all the evaluations using Pytorch 2.1.0 and run them on a single NVIDIA RTX A6000 GPU (48G).
@@
(c) MF-it (CLIP)
(d) MF-ii (CLIP)
(e) MF-it (EVA-CLIP)
(f) MF-ii (EVA-CLIP)
(g) InstructTA-MF-it
(h) InstructTA-MF-ii
Figure 4: Visualization examples of various targeted attack methods on InstructBLIP. “Explain the visual content of the image in great detail.” is a real instruction.
@@ 4.2 Results
Targeted Attack Performance. As illustrated in Table 1, we transfer adversarial samples generated by different targeted attack methods to victim LVLMs and mislead them to produce target responses. For each adversarial example, when attacking a LVLM, the target text and the assigned instruction may be different from other adversarial examples, since we randomly selected 1000 pairs of target texts and instructions from LLaVA-Instruct-150K [27]. Subsequently, we evaluate the similarity between the generated response $\boldsymbol{y}^{\prime}$ of the victim LVLM and the targeted text $\boldsymbol{y}_{t}$ by utilizing diverse forms of CLIP text encoders, where $\boldsymbol{y}^{\prime}=F(\boldsymbol{x}^{\prime},\boldsymbol{p})$. In addition, we count the number of successful targeted adversarial samples, as judged by GPT-4.
@@
Table 1 reports the evaluation results of different targeted attack methods on BLIP-2, InstructBLIP and MiniGPT-4. We compare InstructTA with MF-it and MF-ii, which are the two recent attacks on LVLMs. Overall, we observe that InstructTA consistently outperforms MF-it and MF-ii in terms of CLIP score and the number of successful attacks. In fact, it can be seen that MF-it and MF-ii with CLIP have almost the same results as the clean samples, indicating no targeted attack effect. We believe this is reasonable, because ViT-B/32 in CLIP shares a different visual encoder than that of the target LVLM. In contrast, MF-it (EVA-CLIP), MF-ii (EVA-CLIP) and our methods all notably outperform the targeted attack results on clean samples because they share the same visual encoder as the victim model. Overall, we interpret these results justify the necessity of using the same visual encoder as the victim model.
More importantly, both InstructTA-MF-it and InstructTA-MF-ii are distinctly better than MF-it and MF-ii in terms of CLIP-score and the number of successful attacks. These phenomena demonstrate that our method can effectively improve the transferability of the targeted attack on LVLMs. One reason is that we enhance the robustness of AEs to varying instructions. Another reason is that our dual attacks (InstructTA-MF-it/InstructTA-MF-ii) extracts informative features containing global and instruction-aware knowledges, benefiting to precise targeted attacks.
@@
Table 2: Ablation study of the proposed method on BLIP-2. CLIP score [46] is computed by ensemble CLIP text encoders. Attack success rate is the ratio between the No. of attack successes and the total No. of samples.
@@
Attack method CLIP score ($\uparrow$) Attack success rate ($\uparrow$) InstructTA-MF-it 0.622 0.652 InstructTA-MF-ii 0.616 0.619 InstructTA 0.606 0.624 InstructTA-woG 0.571 0.448
@@
Visualization Examples. Fig. 3 has displayed a sample inferred instruction and its sibling instructions rephrased by ChatGPT. It can be seen that the instructions inferred by GPT-4 are similar to the real instruction. The rephrased instructions are even more similar to the original sentence.
@@
Moreover, we also provide some visual results of different attack methods on attacking InstructBLIP, as shown in Fig. 4. Overall, it can be observed that the generated responses of our methods are closer to the target texts.
@@
Budget $\epsilon$ Text encoder (pretrained) for evaluation Attack success rate RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 2 0.437 0.547 0.480 0.492 0.330 0.457 0.154 4 0.560 0.651 0.602 0.612 0.470 0.579 0.450 8 0.604 0.691 0.644 0.656 0.515 0.622 0.652 16 0.617 0.702 0.657 0.669 0.527 0.634 0.746 64 0.617 0.704 0.658 0.671 0.529 0.636 0.740
Figure 5: To explore the impact of varying $\epsilon$ values within the InstructTA-MF-it, we conducted experiments aiming to achieve different levels of perturbed images on BLIP-2, i.e., referred to as the AE $\boldsymbol{x}^{\prime}$. Our findings indicate a degradation in the visual quality of $\boldsymbol{x}^{\prime}$, as quantified by the LPIPS [44] distance between the original image $\boldsymbol{x}$ and the adversarial image $\boldsymbol{x}^{\prime}$. Simultaneously, the effectiveness of targeted response generation reaches a saturation point. Consequently, it is crucial to establish an appropriate perturbation budget, such as $\epsilon=8$, to effectively balance the image quality and the targeted attack performance.
@@ 4.3 Ablation Study
We also launched ablation study to explore the contribution of major components in the proposed method. At this step, we conducted experiments on BLIP-2 with two key component ablated sequentially: 1) removing MF-it and MF-ii (denoted as InstructTA), and 2) launching optimization without CLIP embedding matching nor ChatGPT for paraphrase (in accordance with Eqn. 4; denoted as InstructTA-woG).
The results are reported in Table 2. It is observed that the performances degrade when one or more parts are removed. In particular, the InstructTA-woG setting exhibits a success rate downgrading of over 31% in comparison to the InstructTA-MF-it setting. Overall, we interpret the ablation study results as reasonable, illustrating that all the constituent components have been thoughtfully devised and exhibit congruity, culminating in an efficacious solution for the targeted attack.
@@ 4.4 Discussion
Effect of $\epsilon$. To explore the effect of the perturbation magnitude $\epsilon$ on targeted attack performance, we study the attack results of InstructTA-MF-it toward BLIP-2 under various $\epsilon$. The results are reported in Table 3. It is evident that with $\epsilon$ increasing, both CLIP scores and attack success rates rise, illustrating stronger attack performance. Nevertheless, it is seen that the attack performance becomes gradually saturated when $\epsilon$ is greater than $16$. Moreover, as shown in Fig. 5, an overly large $\epsilon$ (e.g., 16 and 64) has made the perturbation perceptible to naked eyes. Thus, we suggest to explore a proper $\epsilon$ for the targeted attack, e.g., $\epsilon=8$, such that the attack performance can be effectively improved while the perturbation is imperceptible.
Table 4: Average CLIP score ($\uparrow$) between different versions of instructions. Each type of instruction have 1,000 samples.
Instruction Rephrased instruction $\boldsymbol{p}_{i/j}^{\prime}$ Real instruction $\boldsymbol{p}$ Inferred instruction $\boldsymbol{p}^{\prime}$ 0.942 0.883 Real instruction $\boldsymbol{p}$ 0.882 1.000
Assessment of Generated Instructions. To quantitatively evaluate the usability of the generated instructions, we collect a paraphrased instruction for each inferred instruction used in main experiment and provide average CLIP score values between them, as reported in Table 4. It is evident that there exists a high similarity score between the rephrased and the inferred instructions; moreover, the rephrased and the inferred instructions both show high similarity with the real instructions. We interpret the results as encouraging, demonstrating the high quality of employed instructions.
@@ 5 Conclusion, Ethics, and Mitigation
Our work highlights the vulnerability of LVLMs to adversarial attacks under a practical, gray-box scenario. InstructTA delivers targeted attacks on LVLMs with high transferability, demonstrating the importance of instruction-based techniques and the fusion of visual and language information in boosting adversarial attacks. The aim of our work is to enhance the resilience of LVLMs. Ethically, since we discovered adversarial examples, we minimized harm by (1) avoiding any damage to real users and (2) disclosing the detailed attacking process and methods used to find them in this paper. Overall, we believe that the security of the LVLM ecosystem is best advanced by responsible researchers investigating these problems. Looking ahead, we envision the feasibility of launching various mitigation methods to improve the robustness of LVLMs against adversarial attacks. For instance, we can leverage the adversarial training framework [30] to train LVLMs with adversarial examples. We can also explore the possibility of incorporating adversarial training into the pre-training stage of LVLMs or their visual encoders. Moreover, we can also explore detecting adversarial examples by leveraging the discrepancy between the instruction-aware features of the adversarial examples and the original images.
Supplementary Material
6 Attack Cases
Fig. 6 shows an example to attack BLIP-2 for different methods. It can be seen from the figure that CLIP-based MF-it and MF-ii shows no targeted attack effect since they have the similar results with the clean sample. Interestingly, half of the response (i.e., “a man and a woman”) of the adversarial sample generated by MF-it (EVA-CLIP) matches the original image and half (i.e., “elephants in a zoo”) matches the target image. MF-ii (EVA-CLIP), InstructTA-MF-it and InstructTA-MF-ii demonstrate the targeted attack effect. Importantly, the description of our InstructTA-MF-ii best matches the target image.
7 Results with Complex Reasoning Questions
Inferring from images can demonstrate the model’s deep understanding of the image semantics. The box below shows some reasoning instructions from complex_reasoning_77k in LLaVA-Instruct-150K [27]. 1. How can pedestrians ensure their safety when navigating this busy city scene at night? 2. What might the cat be curious about on the floor? 3. What suggestions can be made for individuals wishing to serve these dough balls in an aesthetically pleasing manner? 4. What might be the man’s intention in the kitchen? 5. How does the position of the players on the field affect their performance? 6. What unique feature can be observed about the desserts in the scene? 7. What type of event might be taking place in this scene? 8. What could be a possible reason for the dog and cat’s interaction? 9. In terms of architectural features, why might this clock tower be studied or admired? 10. What tactic might the man be employing during the tennis match? 11. What could be the significance or purpose of a single pink rose placed in a glass with water? 12. What can you infer about the laptop owner’s personal style or interests? 13. How is the truck being used during this event, and what challenges might it face in this context? 14. Why might the vandalized stop sign on the school bus cause concern? …
To further evaluate the effectiveness of the proposed methods, we provide transfer attack results for complex reasoning instructions, as portrayed in Table 6. All methods, even clean samples, have obvious targeted attack performance. This may be because some of the selected reasoning instructions can be answered without the need for image information, e.g., “Why might the vandalized stop sign on the school bus cause concern?”. Importantly, our methods have better transfer attack performances in terms of both CLIP score and the No. of successes. It implies that InstructTA still performs well even when faced with complex inference prompts.
8 Victim Models
Table 5 shows the architectures of the target LVLMs.
Table 5: The architectures of the victim LVLMs.
LVLM Visual Encoder Projector LLM BLIP-2 ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Flan-T5-XXL InstructBLIP ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Vicuna-7b-v1.1 MiniGPT-4 ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Vicuna-v0-13b
Figure 6: An instance of different targeted attack methods on BLIP-2. The real instruction is “Can you describe the main features of this image for me?”. The target text is “The image depicts a large elephant and a smaller elephant standing in a grassy area next to a stone wall, likely within a zoo enclosure. The two elephants appear to be gathered near a flock of black birds. The birds can be seen scattered around the enclosure, with some standing near the elephants and others in the background. The scene captures the interaction between the elephants and the birds, creating a serene and captivating moment.”
Table 6: Targeted attacks against victim BLIP-2 with reasoning instructions. We randomly select a total of 1,000 clean images from the ImageNet-1K [13] validation set. For each of these benign sample, a target text and a prompt are randomly assigned from complex_reasoning_77k in LLaVA-Instruct-150K [27]. Our metrics involves the computation of the CLIP score (normal-$\uparrow$) [34, 46] between the target texts and the generated responses of AEs on the attacked LVLM, which was performed using several CLIP text encoders and their ensemble. We also report the No. of successful attacks within 1,000 adversarial samples, determined by GPT-4. CLIP and EVA-CLIP in parentheses indicate that they are used as surrogate models for transfer attack.

### v3

@@ 4.1 Evaluation Setup
Datasets. The ImageNet-1K [15] comprises 3 subsets, including training, validation and test sets. We randomly select 1,000 images from the validation set to serve as the benign samples ($\boldsymbol{x}$) for the targeted attack. To establish the instruction-answer pairs as tested instructions and target texts, we adapt 11 instructions (see the supplementary material) from LLaVA-Instruct-150K [32] and employ GPT-4 to generate the corresponding answers. Specifically, we pick 1,000 images from the validation set of COCO [30], and for each image, we randomly select an instruction from these 11 instructions and utilize the GPT-4V [38] API to output the response. In addition, we use Stable Diffusion ($h_{\xi}$) to generate target images from the target texts, following [55].
@@
Protocols. Following [55], we adopt CLIP-score ($\uparrow$) to evaluate the targeted attack performance, which calculates the semantic similarity between the generated response of victim LVLM and the target text using CLIP text encoders. Additionally, we record the number of successful attacks (NoS) and attack success rate (ASR) by using GPT-4 to determine whether two texts have the same meaning. The prompt in this operation is as follows:
Attack success prompt Determine whether these two texts describe the same objects or things, you only need to answer yes or no: 1. $\boldsymbol{y}^{\prime}$ 2. $\boldsymbol{y}_{t}$
where $\boldsymbol{y}^{\prime}$ denotes the output response of the adversarial example on the victim model, and $\boldsymbol{y}_{t}$ is the target response.
@@
Baselines. This paper provides evaluations on five victim LVLMs, including BLIP-2 [28], InstructBLIP [14], MiniGPT-4 [57], LLaVA [32], and CogVLM [46]. In detail, we select flan-t5-xxl [^2] https://huggingface.co/google/flan-t5-xxl, vicuna-7b-v1.1333https://huggingface.co/lmsys/vicuna-7b-v1.1, vicuna-v0-13b [^4] https://huggingface.co/Vision-CAIR/vicuna, LLaMA-2-13B [^5] https://huggingface.co/liuhaotian/llava-llama-2-13b-chat-lightning-preview and vicuna-7b-v1.5666https://huggingface.co/lmsys/vicuna-7b-v1.5 as LLMs for BLIP-2, InstructBLIP, MiniGPT-4, LLaVA and CogVLM [^7] https://huggingface.co/THUDM/cogvlm-base-224-hf, respectively. For BLIP-2, InstructBLIP and MiniGPT-4, their visual encoders are all ViT-G/14 [51] borrowed from EVA-CLIP [19]. The vision encoders for LLaVA and CogVLM are built on ViT-L/14 [40] and EVA2-CLIP-E [43], respectively. To evaluate the targeted attack performance, InstructTA is compared with existing baselines, i.e., MF-tt, MF-it and MF-ii [55]. MF-tt is a query-based black-box attack method that only accesses the input and output of the target model.
@@
Implementation Details. All images are resized to $224\times 224$ before feeding in substitute models and LVLMs. MF-it and MF-ii choose the same vision encoders as the victim LVLMs as surrogate models for targeted attacks. For InstructTA, we choose the vision encoder (ViT-G/14) and Q-Former from InstructBLIP as the surrogate model. In other words, ViT-G/14 and Q-Former are the visual encoder and projector in Fig. 1, respectively. For BLIP-2, InstructBLIP and MiniGPT-4, our surrogate models of InstructTA share the same vision encoders as the target models, regardless of the instruction-aware surrogate model or MF-it used in the dual targeted attack. For LLaVA and CogVLM, only MF-it in InstructTA (Eqn. 5) keeps the same vision encoder as the victim model. For PGD, $\eta$ and $T$ are set to 1 and 100, respectively. The maximum perturbation magnitude $\epsilon$ is set to $8$. The settings of MF-tt, MF-it and MF-ii are all consistent with the original literature [55]. We implement all the evaluations using Pytorch 2.1.0 and run them on a single NVIDIA RTX A6000 GPU (48G).
Table 1: Targeted attacks against victim LVLMs. A total of 1,000 clean images denoted as $\boldsymbol{x}$ are sampled from the ImageNet-1K [15] validation set. For each of these benign images, a targeted text $\boldsymbol{y}_{t}$ and a prompt $\boldsymbol{p}$ are randomly assigned from our generated instruction-answer paired set (see § 4.1). Our analysis involves the computation of the CLIP-score ($\uparrow$) [40, 55] between the target texts and the generated responses of AEs on attacked LVLMs, which was performed using several CLIP text encoders and their ensemble. We also report the No. of successful attacks (NoS) within 1,000 adversarial samples, determined by GPT-4.
LVLM Attack method Text encoder (pretrained) for evaluation NoS ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.281 0.395 0.323 0.333 0.184 0.303 1 MF-tt [55] 0.284 0.400 0.326 0.337 0.185 0.306 0 MF-it [55] 0.533 0.634 0.575 0.581 0.469 0.559 401 MF-ii [55] 0.536 0.634 0.575 0.583 0.471 0.560 414 InstructTA 0.562 0.660 0.604 0.611 0.502 0.588 519 InstructBLIP Clean image 0.329 0.450 0.358 0.361 0.246 0.349 1 MF-tt [55] 0.332 0.452 0.361 0.363 0.249 0.351 1 MF-it [55] 0.650 0.714 0.675 0.676 0.608 0.665 233 MF-ii [55] 0.638 0.707 0.666 0.667 0.594 0.654 223 InstructTA 0.683 0.740 0.709 0.708 0.645 0.697 300 MiniGPT-4 Clean image 0.324 0.448 0.348 0.351 0.244 0.343 2 MF-tt [55] 0.332 0.455 0.353 0.356 0.243 0.348 0 MF-it [55] 0.593 0.672 0.611 0.618 0.538 0.606 121 MF-ii [55] 0.602 0.682 0.621 0.623 0.548 0.615 128 InstructTA 0.633 0.701 0.650 0.654 0.578 0.643 164 LLaVA Clean image 0.325 0.465 0.374 0.383 0.216 0.352 1 MF-tt [55] 0.329 0.466 0.376 0.386 0.217 0.355 1 MF-it [55] 0.368 0.502 0.418 0.424 0.268 0.396 32 MF-ii [55] 0.360 0.496 0.407 0.415 0.256 0.387 30 InstructTA 0.395 0.524 0.445 0.449 0.300 0.423 36 CogVLM Clean image 0.301 0.425 0.319 0.334 0.208 0.318 4 MF-tt [55] 0.307 0.432 0.327 0.342 0.212 0.324 8 MF-it [55] 0.493 0.597 0.506 0.513 0.417 0.505 71 MF-ii [55] 0.499 0.600 0.516 0.521 0.423 0.512 63 InstructTA 0.528 0.624 0.539 0.545 0.452 0.537 103
@@
(c) MF-tt
(d) MF-it
(e) MF-ii
(f) InstructTA
Figure 4: Visualization examples of various targeted attack methods on InstructBLIP. “What are the essential components depicted in this image?" is a real instruction.
@@ 4.2 Results
Targeted Attack Performance. As illustrated in Table 1, we transfer adversarial samples generated by different targeted attack methods to victim LVLMs and mislead them to produce target responses. For each adversarial example, when attacking an LVLM, the target text and the assigned instruction may be different from other adversarial examples, since we randomly selected 1000 pairs of target texts and instructions from our generated instruction-answer paired set (see § 4.1). Subsequently, we evaluate the similarity between the generated response $\boldsymbol{y}^{\prime}$ of the victim LVLM and the targeted text $\boldsymbol{y}_{t}$ by utilizing diverse forms of CLIP text encoders, where $\boldsymbol{y}^{\prime}=F(\boldsymbol{x}^{\prime},\boldsymbol{p})$. In addition, we tabulate the count of successful targeted adversarial samples (NoS), as evaluated by GPT-4.
@@
Table 1 reports the evaluation results of different targeted attack methods on BLIP-2, InstructBLIP and MiniGPT-4. We compare InstructTA with MF-tt, MF-it and MF-ii, which are the three recent attacks on LVLMs. Overall, we observe that InstructTA consistently outperforms MF-tt, MF-it and MF-ii in terms of CLIP score and NoS. In fact, it can be seen that MF-tt has almost the same results as the clean samples, indicating no promising targeted attack effect. In contrast, MF-it, MF-ii, and our method exhibit significant improvements over the clean samples on targeted attack results. This observation indicates the effectiveness of targeted attack methods against victim LVLMs.
Significantly, InstructTA surpasses MF-it and MF-ii in terms of CLIP-score and NoS, underscoring its superiority. These observations highlight the capacity of our method to enhance the transferability of targeted attacks on LVLMs. This is attributable to two factors. Firstly, we enhance the resilience of adversarial examples to diverse instructions, thereby bolstering their robustness. Secondly, our dual attack scheme extracts informative features that encompass both global and instruction-aware knowledge, thus facilitating precise targeted attacks.
@@
Table 2: Ablation study of the proposed method on BLIP-2. CLIP-score [55] is computed by ensemble CLIP text encoders. Attack success rate (ASR) is the ratio between the No. of attack successes and the total No. of samples.
@@
         Attack method          CLIP-score ($\uparrow$)          ASR ($\uparrow$)          InstructTA          0.588          0.519          InstructTA-woMF          0.574          0.472          InstructTA-woMFG          0.521          0.324
Generalization of InstructTA. We argue that InstructTA exhibits a degree of generalization in attacking LVLMs, even when the vision encoder of the instruction-aware surrogate model differs from that of the target LVLM. As evidenced in Table 1, the surrogate model employed is ViT-G/14 from EVA-CLIP [19], distinct from the ViT-G/14 and EVA02-bigE-14 utilized in LLaVA and CogVLM respectively. Despite this discrepancy, we observe that InstructTA outperforms previous methods in terms of targeted attack effectiveness. Hence, InstructTA can enhance its attack transferability beyond MF-it without necessitating the maintenance of an identical visual encoder to the victim LVLM.
@@
Visualization Examples. Fig. 3 has displayed a sample inferred instruction and its sibling instructions rephrased by GPT-4. It can be seen that the instruction inferred by GPT-4 is similar to the real instruction. The rephrased instructions are even more similar to the original sentence.
@@
Moreover, we also provide some visual results of different attack methods on attacking InstructBLIP, as shown in Fig. 4. Overall, it can be observed that the generated response of our methods is closer to the target text.
@@
Budget $\epsilon$ Text encoder (pretrained) for evaluation ASR ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 2 0.406 0.516 0.451 0.458 0.326 0.431 0.127 4 0.510 0.613 0.552 0.558 0.445 0.535 0.349 8 0.562 0.660 0.604 0.611 0.502 0.588 0.519 16 0.570 0.669 0.610 0.617 0.509 0.595 0.605 64 0.575 0.673 0.616 0.623 0.516 0.600 0.604
Figure 5: To explore the impact of varying $\epsilon$ values within the InstructTA, we conducted experiments aiming to achieve different levels of perturbed images on BLIP-2, i.e., referred to as the AE $\boldsymbol{x}^{\prime}$. Our findings indicate a degradation in the visual quality of $\boldsymbol{x}^{\prime}$, as quantified by the LPIPS [53] distance between the original image $\boldsymbol{x}$ and the adversarial image $\boldsymbol{x}^{\prime}$. Simultaneously, the effectiveness of targeted response generation reaches a saturation point. Consequently, it is crucial to establish an appropriate perturbation budget, such as $\epsilon=8$, to effectively balance the image quality and the targeted attack performance.
@@ 4.3 Ablation Study
We also launched an ablation study to explore the contribution of major components in the proposed method. At this step, we conducted experiments on BLIP-2 with two key components ablated sequentially: 1) removing MF-it (denoted as InstructTA-woMF), and 2) launching optimization without MF-it nor GPT-4 for paraphrases (in accordance with Eqn. 3; denoted as InstructTA-woMFG).
The results are reported in Table 2. It is observed that the performances degrade when one or more parts are removed. In particular, the InstructTA-woMFG setting exhibits a success rate downgrading of over 19% in comparison to the complete InstructTA framework. Overall, we interpret the ablation study results as reasonable, illustrating that all the constituent components have been thoughtfully devised and exhibit congruity, culminating in an efficacious solution for the targeted attack.
@@ 4.4 Discussion
Effect of $\epsilon$. To explore the effect of the perturbation magnitude $\epsilon$ on targeted attack performance, we study the attack results of InstructTA toward BLIP-2 under various $\epsilon$. The results are reported in Table 3. It is evident that with $\epsilon$ increasing, both CLIP-scores and attack success rates rise, illustrating stronger attack performance. Nevertheless, it is seen that the attack performance becomes gradually saturated when $\epsilon$ is greater than $16$. Moreover, as shown in Fig. 5, an overly large $\epsilon$ (e.g., 16 and 64) has made the perturbation perceptible to naked eyes. Thus, we suggest to explore a proper $\epsilon$ for the targeted attack, e.g., $\epsilon=8$, such that the attack performance can be effectively improved while the perturbation is imperceptible.
Effect of count $n$. In rephrasing prompt $\boldsymbol{p}_{r}$, we use $n-1$ to control the number of instructions rephrased by GPT-4. With the original instruction, we totally collect $n$ instructions to participate in the tuning for our targeted attack. To explore the effect of $n$ on targeted attack performance, we perform InstructTA on BLIP-2 with varying $n$. The results are illustrated in Table 4. We can observe that CLIP scores are roughly the same for different $n$. It is worth noting that as n increases, the attack success rate (ASR) also increases. Therefore, increasing the number of paraphrased instructions can further improve ASR.
Table 4: Targeted attack performance for different $n$.
Parameter $n$ Text encoder (pretrained) for evaluation ASR ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 1 0.560 0.658 0.600 0.606 0.498 0.584 0.504 5 0.560 0.657 0.600 0.606 0.498 0.584 0.504 10 0.562 0.660 0.604 0.611 0.502 0.588 0.519 50 0.562 0.659 0.601 0.607 0.499 0.586 0.529
Table 5: Average CLIP-score ($\uparrow$) between different versions of instructions. Each type of instruction has 1,000 samples.
Instruction Rephrased instruction $\boldsymbol{p}_{i/j}^{\prime}$ Real instruction $\boldsymbol{p}$ Inferred instruction $\boldsymbol{p}^{\prime}$ 0.943 0.829 Real instruction $\boldsymbol{p}$ 0.828 1.000
Assessment of Generated Instructions. To quantitatively evaluate the usability of the generated instructions, we collect a paraphrased instruction for each inferred instruction used in the main experiment and provide average CLIP-score values between them, as reported in Table 5. It is evident that there exists a high similarity score between the rephrased and the inferred instructions; moreover, the rephrased and the inferred instructions both show high similarity with the real instructions. We interpret the results as encouraging, demonstrating the high quality of employed instructions.
@@ 5 Conclusion, Ethics, and Mitigation
Our work highlights the vulnerability of LVLMs to adversarial attacks under a practical, gray-box scenario. InstructTA delivers targeted attacks on LVLMs with high transferability, demonstrating the importance of instruction-based techniques and the fusion of visual and language information in boosting adversarial attacks. The aim of our work is to enhance the resilience of LVLMs. Ethically, since we discovered adversarial examples, we minimized harm by (1) avoiding any damage to real users and (2) disclosing the detailed attacking process and methods used to find them in this paper. Overall, we believe that the security of the LVLM ecosystem is best advanced by responsible researchers investigating these problems. Looking ahead, we envision the feasibility of launching various mitigation methods to improve the robustness of LVLMs against adversarial attacks. For instance, we can leverage the adversarial training framework [35] to train LVLMs with adversarial examples. We can also explore the possibility of incorporating adversarial training into the pre-training stage of LVLMs or their vision encoders. Moreover, we can also explore detecting adversarial examples by leveraging the discrepancy between the instruction-aware features of the adversarial examples and the original images.
Appendix 0.A Real Instructions
Here are the 11 real instructions adapted from LLaVA-Instruct-150K:
  • What’s your interpretation of what’s happening in this picture?
  • What is taking place in this scenario?
  • Describe the image’s visual elements with extensive detail.
  • Could you explain the key elements of this picture to me?
  • What are the essential components depicted in this image?
  • What does this picture represent?
  • Could you provide more detail on the components of the image shown?
  • Compose an elaborate depiction of the presented picture.
  • Examine the picture thoroughly and with attention to detail.
  • Provide a description of the image below.
  • What can you observe occurring in this picture?
Table A6: Targeted attack performance with shuffled instructions.